OpenStudy (anonymous):

i need help with dividing polynomials by monomials 9x^14-81x^8/27x^12

OpenStudy (saifoo.khan):

\[\Large \frac{9x^{14}-81x^8}{27x^{12 }}\to \frac{9x^8(x^6-9)}{27x^{12}}\to Solve\]

OpenStudy (anonymous):

(a^b) / (a^c) = a^(b-c) = 1 / a^(c-b)